### v1

# On irreducibility of six-dimensional compatible systems of $\mathbb{Q}$

Abstract.
We study irreducibility of 6-dimensional strictly compatible systems of $\mathbb{Q}$ with distinct Hodge-Tate weights. We prove if one of the representations is Lie-irreducible, then all but finitely many of them are (Lie-)irreducible. If one of them is irreducible but not Lie-irreducible, we prove the same result under certain self-duality condition.
1. Introduction
Given a rational prime number $\ell$ and an elliptic curve $E$ over a number field $K$. The $\ell$-adic Tate module $\rho_{\ell}:=T_{\ell}(E)$ and $\ell$-torsion points $\overline{\rho}_{\ell}:=E[\ell]$ form classical examples of (2-dimensional) compatible systems and mod $\ell$ compatible systems. The following result asserts that they have uniform type of irreducible decompositions. Here (i) can be regarded as a weak version of Serre’s famous big image result in [Se72].
Theorem 1.1.
Given an elliptic curve $E$ over a number field $K$. Consider compatible system $\{\rho_{\ell}\}$ and mod $\ell$ compatible system $\{\overline{\rho}_{\ell}\}$ of $\mathrm{Gal}_{K}$.
  1. (i)
    If $E$ has no complex multiplication over $\overline{K}$, then each $\rho_{\ell}$ is absolutely Lie-irreducible and $\overline{\rho_{\ell}}$ is absolutely irreducible for sufficiently large $\ell$.
  2. (ii)
    If $E$ has complex multiplication (over $K$), then after possibly enlarging the coefficients, $\{\rho_{\ell}\}$ can be written as a direct sum of two 1-dimensional compatible systems.
  3. (iii)
    If $E$ has potential complex multiplication but not over $K$, then $\{\rho_{\ell}\}$ is absolutely irreducible and induced by a 1-dimensional compatible system over a quadratic extension of $K$ after necessarily enlarging the coefficients. Also $\overline{\rho_{\ell}}$ is absolutely irreducible for sufficiently large $\ell$.
It is generally believed the above patterns are true for arbitrary semisimple compatible systems. More precisely we have the following 1.2. Note that in the above elliptic curve cases, the compatible systems are regular.
Conjecture 1.2.
Given an $n$-dimensional strictly compatible system $\mathcal{M}$ of a number field $K$.
  1. (i)
    $\mathcal{M}$ can be decomposed as the direct sum of irreducible strictly compatible systems.
  2. (ii)
    For irreducible $\mathcal{M}$, the representations are residually irreducible except for finitely many of them.
  3. (iii)
    For irreducible and regular $\mathcal{M}$, it can be written as induced representations of a Lie-irreducible strictly compatible system.
For a summary of history of low dimensional cases, see [Hu23a, Section 1.2]. For $K=\mathbb{Q}$ and under regularity condition, case $n=4$ was treated in [Hu23a]; case $n=5$ (includes partial irregular cases) was treated in [DWW24]. The present paper treats $n=6$. The main results are:
Theorem 1.3.
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational regular strictly compatible system of $\mathbb{Q}$. Suppose some $\rho_{\lambda_{0}}$ is irreducible.
  1. (i)
    If $\rho_{\lambda_{0}}$ is Lie-irreducible, then $\rho_{\lambda}$ is (Lie-)irreducible for all but finitely many $\lambda$.
  2. (ii)
    If $\rho_{\lambda_{0}}$ is not Lie-irreducible but satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ with some character $\chi$, then $\rho_{\lambda}$ is irreducible for all but finitely many $\lambda$.
  3. (iii)
    If $\rho_{\lambda_{0}}$ satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ with some odd character $\chi$, then the semisimple reduction $\overline{\rho_{\lambda}}^{\text{ss}}$ is irreducible for all but finitely many $\lambda$.
Corollary 1.4.
Let $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational pure essentially self-dual and totally odd regular strictly compatible system of a totally real field $K$. Then there exists $r\in\mathbb{N}$ such that:
  1. (i)
    For all but finitely many $\lambda$, one has irreducible decompositions:
    $\rho_{\lambda}=\sigma_{\lambda,1}\oplus\sigma_{\lambda,2}\oplus\cdots\oplus\sigma_{\lambda,r}$
  2. (ii)
    Each $\{\sigma_{\lambda,i}\}$ can be extended to a strictly compatible system.
We arrange the article as following. In section 2 we give necessary preliminaries for the proofs, including certain $\ell$-independence properties, big image results, potential automorphy theorem which is the main tool of the proof, and certain results on the semisimple reductions. In section 3 we prove the main results.
2. Preliminaries
2.1. Compatible systems
Definition 2.1.
Let $K$ be a number field. An $n$-dimensional $E$-rational strictly compatible system of $\mathrm{Gal}_{K}$ is the datum
$\mathcal{M}=\left(E,S,\{P_{v}(T)\},\{\rho_{\lambda}\},\{\text{HT}_{\tau}\},\{\text{WD}_{v}\}\right)$
where:
  • $E$ is a number field.
  • $S$ is a finite set of primes in $K$ called exceptional set.
  • $P_{v}(T)\in E[T]$ is a degree $n$ monic polynomial for each prime $v\not\in S$ of $K$.
  • $\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{E_{\lambda}})$ is an $n$-dimensional continuous semisimple Galois representation.
  • $\text{HT}_{\tau}$ is a multiset of $n$ integers for each embedding $\tau:K\hookrightarrow\overline{E}$.
  • $\text{WD}_{v}$ is a semisimple Weil-Deligne representation of $K_{v}$ for each prime $v$.
such that:
  1. (i)
    Each $\rho_{\lambda}$ is a geometric representation in the sense of Fontaine-Mazur with exceptional set $S$, this means
    • $\rho_{\lambda}$ is unramified outside $S\cup S_{\lambda}$, where $S_{\lambda}$ is the primes of $K$ dividing the same rational prime as $\lambda$.
    • If $v\in S_{\lambda}$ then $\rho_{\lambda}|_{\mathrm{Gal}_{K_{v}}}$ is de Rham, and moreover is crystalline if $v\not\in S$.
  2. (ii)
    For each $v\not\in S\cup S_{\lambda}$, the characteristic polynomial of $\rho_{\lambda}(\text{Frob}_{v})$ is $P_{v}(T)$.
  3. (iii)
    For each embedding $\tau:K\hookrightarrow\overline{E}$ and each $E$-embedding $\overline{E}\hookrightarrow\overline{E_{\lambda}}$, the Hodge-Tate weights of $\rho_{\lambda}$ is $\text{HT}_{\tau}$.
  4. (iv)
    For each $v\not\in S_{\lambda}$ and each isomorphism $\iota:\overline{E_{\lambda}}\cong\mathbb{C}$, the Frobenius semisimplified Weil-Deligne representation $\iota\text{WD}(\rho_{\lambda}|_{\mathrm{Gal}_{K_{v}}})^{F-\text{ss}}$ is isomorphic to $\text{WD}_{v}$.
Definition 2.2.
A strictly compatible system $\{\rho_{\lambda}\}$ is called regular, if for any embedding $\tau:K\hookrightarrow\overline{E}$, the $\tau$-Hodge-Tate weights are distinct.
Under regularity condition, one can descend the coefficients of a strictly compatible system to $E_{\lambda}$ after possibly enlarging $E$:
Lemma 2.3.
[BLGGT14, Lemma 5.3.1.(3)] Let $\{\rho_{\lambda}\}$ be an $E$-rational strictly compatible system of $K$. Suppose $\mathcal{M}$ is regular, then after replacing $E$ with a finite extension, we may assume that for any open subgroup $H$ of $\mathrm{Gal}_{K}$, any $\lambda$ and any $H$-subrepresentation $\sigma$ of $\rho_{\lambda}$, the representation $\sigma$ is defined over $E_{\lambda}$.
Definition 2.4.
A strictly compatible system $\{\rho_{\lambda}\}$ is called pure of weight $\omega$, if for each $v\not\in S$, for each root $\alpha$ of $P_{v}(T)$ in $\overline{E}$ and for each embedding $i:\overline{E}\hookrightarrow\mathbb{C}$, one has $|i\alpha|^{2}=(\#\kappa(v))^{\omega}$, where $\kappa(v)$ is the residue field at $v$.
Certain results we use work well under weaker conditions than 2.1. We mention this definition for accuracy.
Definition 2.5.
Let $K$ be a number field. An $n$-dimensional $E$-rational Serre compatible system of $\mathrm{Gal}_{K}$ is the datum
$\mathcal{M}=\left(E,S,\{p_{v}(T)\},\{\rho_{\lambda}\}\right)$
where:
  • $E$ is a number field.
  • $S$ is a finite set of primes in $K$ called exceptional set.
  • $P_{v}(T)\in E[T]$ is a degree $n$ monic polynomial for each prime $v\not\in S$ of $K$.
  • $\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{n}(E_{\lambda})$ is an $n$-dimensional continuous semisimple Galois representation.
such that:
  1. (i)
    $\rho_{\lambda}$ is unramified outside $S\cup S_{\lambda}$, where $S_{\lambda}$ is the primes of $K$ dividing the same rational prime as $\lambda$.
  2. (ii)
    For each $v\not\in S\cup S_{\lambda}$, the characteristic polynomial of $\rho_{\lambda}(\text{Frob}_{v})$ is $P_{v}(T)$.
2.2. $\lambda$-independence
Definition 2.6.
Given a $\lambda$-adic semisimple continuous Galois representation
$\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{E_{\lambda}})$
Its algebraic monodromy group $\textbf{G}_{\lambda}$ is defined to be the Zariski closure of the image inside algebraic group $\text{GL}_{n,\overline{E_{\lambda}}}$.
Lemma 2.7.
Given a $\lambda$-adic Galois representation $\rho$. Assume $\sigma$ is a subquotient of $\rho$. Then the algebraic monodromy groups of $\rho$ and $\rho\oplus\sigma$ coincide.
Proof.
The algebraic monodromy group of $\rho$ is the fundamental group of the Tannakian category generated by $\rho$ inside the category of all $\lambda$-adic Galois representations. The result follows from the fact $\rho$ and $\rho\oplus\sigma$ generate the same Tannakian category. ∎
Due to semisimplicity, the identity component $\textbf{G}_{\lambda}^{\circ}$ is a reductive group. We write $\textbf{G}_{\lambda}^{\text{der}}$ to be the derived subgroup of $\textbf{G}_{\lambda}^{\circ}$, which is semisimple.
Definition 2.8.
Let $F$ be a field and $\textbf{G}\subseteq\mathrm{GL}_{n,F}$ a reductive subgroup.
  1. (i)
    Denote by T the maximal torus of $\textbf{G}\times\overline{F}$. Then the formal character of G is the conjugacy class of T in $\mathrm{GL}_{n,\overline{F}}$.
  2. (ii)
    Dnote by $\textbf{T}^{\prime}$ the maximal torus of $\textbf{G}^{\mathrm{der}}\times\overline{F}$. Then the formal bi-character of G is the conjugacy class of the chain $^{\prime}\subseteq\textbf{T}$ T in $\mathrm{GL}_{n,\overline{F}}$.
  3. (iii)
    Let $\{F_{i}\}$ be a family of fields and $\{\textbf{G}_{i}\subseteq\mathrm{GL}_{n,F_{i}}\}$ be a family of reductive groups. We say they have same formal bi-character if there exists a chain of subtorus $\textbf{T}_{\mathbb{Z}}^{\prime}\subseteq\textbf{T}_{\mathbb{Z}}\subseteq\mathrm{GL}_{n,\mathbb{Z}}$ such that the chain $\textbf{T}_{\mathbb{Z}}^{\prime}\times\overline{F}_{i}\subseteq\textbf{T}_{\mathbb{Z}}\times\overline{F}_{i}$ is contained in formal bi-character of $\textbf{G}_{i}$ for each $i$.
We have the following standard $\lambda$-independence results on algebraic monodromy groups.
Theorem 2.9.
[Se81], [Se84], [Hu13, Theorem 3.19]. Given an $E$-rational Serre compatible system $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{n}(E_{\lambda})\}$. Denote by $\textbf{G}_{\lambda}$ the algebraic monodromy group of $\rho_{\lambda}\otimes\overline{E_{\lambda}}$
  1. (i)
    The component group $\pi_{0}(\textbf{G}_{\lambda})=\textbf{G}_{\lambda}/\textbf{G}_{\lambda}^{\circ}$ is independent of $\lambda$. In particular the connectedness of $\textbf{G}_{\lambda}$ is independent of $\lambda$.
  2. (ii)
    The formal bi-character of the tautological representation $\textbf{G}_{\lambda}\hookrightarrow\text{GL}_{n,\overline{E_{\lambda}}}$ and hence the rank and semisimple rank of $\textbf{G}_{\lambda}$ are independent of $\lambda$.
Denote by $(\overline{\rho}^{\text{ss}},\overline{V}^{\text{ss}})$ the semisimple reduction of a $\lambda$-adic Galois representation $(\rho,V)$. For the purpose of investigating residual irreducibility, we need the following concept.
Definition/Theorem 2.10.
[Hu23b, Theorem 3.1], [Hu23a, Proposition 2.11]. Given an $n$-dimensional regular $E$-rational strictly compatible system $\{(\rho_{\lambda},V_{\lambda})\}$ of number field $K$. By 2.3, after enlarging $E$, one writes the system as $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(E_{\lambda})\}$. Write $d=[E:\mathbb{Q}]$. By restriction of scalars, we have an $nd$-dimensional $\mathbb{Q}$-rational compatible system:
$\left\{\rho_{\ell}:=\bigoplus_{\lambda|\ell}\rho_{\lambda}:\mathrm{Gal}_{K}\to\left(\mathrm{Res}_{E/\mathbb{Q}}\right)\left(\mathbb{Q}_{\ell}\right)\subseteq\mathrm{GL}_{nd}(\mathbb{Q}_{\ell})\right\}_{\ell}$
  1. (i)
    There exists a finite Galois extension $L/K$ such that, up to isomorphism there exists a unique connected reductive group
    $\underline{G}_{\ell}\subseteq\text{GL}_{nd,\mathbb{F}_{\ell}}$
    for each sufficiently large $\ell$ called algebraic envelope, such that:
    • $\overline{\rho_{\ell}}^{\text{ss}}(\mathrm{Gal}_{L})$ is a subgroup of $\underline{G}_{\ell}(\mathbb{F}_{\ell})$ with index uniformly bounded when $\ell$ varies.
    • $\underline{G}_{\ell}$ acts on the ambient space semisimply.
    • The formal characters of $\underline{G}_{\ell}\hookrightarrow\text{GL}_{nd,\mathbb{F}}$ for all $\lambda$ are bounded.
  2. (ii)
    For all but finitely many $\lambda$, let $\ell$ be the rational prime divided by $\lambda$ and $(\sigma,W)$ be a subrepresentation of $\rho_{\lambda}\otimes\overline{\mathbb{Q}}_{\ell}$. Denote by $\underline{G}_{W}$ the image of $\underline{G}_{\ell}$ in $\text{GL}_{\overline{W}^{\text{ss}}}$, which is called algebraic envelope of $W$.
Theorem 2.11.
[Hu23b, Theorem 3.12],[Hu23a, Theorem 2.10]. Given an $n$-dimensional regular $E$-rational strictly compatible system $\{\rho_{\lambda}\}$ of number field $K$. Except for finitely many $\lambda$, for any subrepresentation $(\sigma,W)$ of $\rho_{\lambda}$ one has:
  1. (i)
    The algebraic envelope $\underline{G}_{W}$ and algebraic monodromy $\textbf{G}_{W}$ of $\sigma$ have the same formal bi-characters.
  2. (ii)
    There exists a finite Galois extension $L/K$, independent of $W$, such that the commutants of $\overline{\sigma}_{\lambda}^{\text{ss}}(\mathrm{Gal}_{L})$ and $\underline{G}_{W}$ (resp. $[\overline{\sigma}_{\lambda}^{\text{ss}}(\mathrm{Gal}_{L}),\overline{\sigma}_{\lambda}^{\text{ss}}(\mathrm{Gal}_{L})]$ and $\underline{G}_{W}^{\text{ss}}$) in $\text{End}(\overline{W})^{\text{ss}}$ are equal. In particular, $\overline{\sigma}_{\lambda}^{\text{ss}}(\mathrm{Gal}_{L})$ (resp. $[\overline{\sigma}_{\lambda}^{\text{ss}}(\mathrm{Gal}_{L}),\overline{\sigma}_{\lambda}^{\text{ss}}(\mathrm{Gal}_{L})]$) is irreducible on $\overline{W}^{\text{ss}}$ if and only if $\underline{G}_{W}$ (resp. $\underline{G}_{W}^{\text{der}}$) is irreducible on $\overline{W}^{\text{ss}}$.
  3. (iii)
    If $\textbf{G}_{W}$ is of type A and $\textbf{G}_{W}^{\circ}\to\mathrm{GL}_{W}$ is irreducible (in particular for Lie-irreducible dimension $\leq 3$ ones), then $\underline{G}_{W}$ and thus $\mathrm{Gal}_{K}$ (resp. $\mathrm{Gal}_{K^{\text{ab}}}$) are irreducible on $\overline{W}^{\mathrm{ss}}$.
  4. (iv)
    If $\sigma$ is irreducible and of type A, then it is residually irreducible.
We mention a recent result on $\ell$-independence of component groups of algebraic envelope groups which is the analogues of Serre’s result 2.9(i). Given a smooth projective variety $X$ over some number field $K$, then the semisimplification of $\ell$-adic cohomology
[…]
(after extending scalars to $\overline{\mathbb{Q}}_{\ell}$) form a $\mathbb{Q}$-rational strictly compatible system of $K$. Such $\{V_{\ell}\}$ is called a compatible system arising from varieties. Define the full algebraic envelope of $V_{\ell}$ when $\ell$ sufficiently large to be:
$\widehat{\underline{G}_{\ell}}:=\text{Im}\overline{\rho_{\ell}}^{ss}\cdot\underline{G}_{\ell}$
Theorem 2.12.
[DH24, Theorem 1.3]. Let $\{\rho_{\ell}:\mathrm{Gal}_{K}\to\text{GL}_{n}(\mathbb{Q}_{\ell})\}_{\ell}$ be a semisimple compatible system that is arising from some variety, with algebraic monodromy groups $\{\textbf{G}_{\ell}\}_{\ell}$ and full algebraic envelopes $\{\widehat{\underline{G}}_{\ell}\}_{\ell\gg 0}$. Let $K^{\text{conn}}/K$ be the finite Galois extension corresponding to $\textbf{G}_{\ell}/\textbf{G}_{\ell}^{\circ}$ which is independent of $\ell$. For all sufficiently large $\ell$, the finite Galois extension corresponding to the morphism
$\mathrm{Gal}_{K}\stackrel{{\scriptstyle\bar{\rho}_{\ell}^{\operatorname{ss}}}}{{\longrightarrow}}\widehat{\underline{G}}_{\ell}(\mathbb{F}_{\ell})\to\widehat{\underline{G}}_{\ell}(\mathbb{F}_{\ell})/\underline{G}_{\ell}(\mathbb{F}_{\ell})$
is $K^{\text{conn}}/K$. In particular, the component groups $\pi_{0}(\textbf{G}_{\ell})=\textbf{G}_{\ell}/\textbf{G}_{\ell}^{\circ}$ and $\pi_{0}(\widehat{\underline{G}}_{\ell})=\widehat{\underline{G}}_{\ell}/\underline{G}_{\ell}$ are naturally isomorphic for all $\ell\gg 0$.
2.3. Lie-irreduciblility
Definition 2.13.
A $\lambda$-adic continuous Galois representation $\rho:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{E_{\lambda}})$ is called Lie-irreducible if for any finite extension $L/K$, the restriction $\rho|_{\mathrm{Gal}_{L}}$ is irreducible.
Given a finite extension of fields $L/K$ and a Galois representation $\sigma$ of $L$, we write the induced representation as
$\text{Ind}_{L}^{K}\sigma:=\text{Ind}_{\mathrm{Gal}_{L}}^{\mathrm{Gal}_{K}}\sigma$
Proposition 2.14.
[Pa19, Proposition 3.4.1, Lemma 3.4.6]. Given an irreducible Hodge-Tate $\lambda$-adic Galois representation $\rho$ of some number field $K$, assume the Hodge-Tate weights are distinct, then either $\rho$ is Lie-irreducible, or is induced from a Lie-irreducible representation $\sigma$ of some finite extension $L/K$:
$\rho=\mathrm{Ind}_{L}^{K}\sigma$
Hence a Hodge-Tate regular semisimple Galois representation $\rho$ can be uniquely written as
$\rho=\oplus_{i}\mathrm{Ind}_{K_{i}}^{K}\sigma_{i}$
where each $\sigma_{i}$ is Lie-irreducible of $K_{i}$. Denote by $\text{Spl}(\rho)$ the set of primes in $K$ which have a split factor in at least one $K_{i}$. Then the usual formula for the trace of an induced representation shows $\text{Spl}(\rho)$ is contained in the set of trace zero (under $\rho$) primes of $K$. [Pa19, Proposition 3.4.9.(1)] shows the Dirichlet density of the two sets coincide. Hence by compatibility of the system we obtain:
Proposition 2.15.
Given a Hodge-Tate regular $E$-rational Serre compatible system $\{\rho_{\lambda}\}$ of some number field $K$. For each $\lambda$ consider the unique decomposition guarateed by 2.14:
$\rho_{\lambda}=\oplus_{i}\mathrm{Ind}_{K_{\lambda,i}}^{K}\sigma_{\lambda,i}$
with each $\sigma_{i}$ being Lie-irreducible of $K_{\lambda,i}$. Denote by $d(\rho_{\lambda})$ the Dirichlet density of primes in $K$ which have a split factor in at least one $K_{\lambda,i}$. Then this density is independent of $\lambda$.
Corollary 2.16.
In particular, under above conditions, if some $\rho_{\lambda_{0}}$ is induced by some representation of a nontrivial field extension of $K$, then each irreducible component of each $\rho_{\lambda}$ cannot be Lie-irreducible.
2.4. Essential self-duality
Definition 2.17.
Let $K$ be a totally real field. Let $E$ be a number field and $\lambda$ be a prime of $E$. A $\lambda$-adic Galois representation $\rho:\text{Gal}_{K}\to\text{GL}_{n}(\overline{E}_{\lambda})$ is called essentially self-dual, if it either factors through $\text{GSp}_{n}(\overline{E}_{\lambda})$ or $\text{GO}_{n}(\overline{E}_{\lambda})$. In particular there exists some continuous character $\chi:\text{Gal}_{K}\to\overline{E}_{\lambda}^{*}$ called similitude character, such that $\rho\cong\rho^{\vee}\otimes\chi$.
Our condition $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ in 1.3.(ii) is slightly general than essential self-duality. If a representation satisfies this condition, again we call $\chi$ the similitude character. Note that the similitude character may not be unique.
Definition 2.18.
A Galois character $\chi$ of a totally real field is called totally odd (respectively, totally even), if for any complex conjugation $c$ one has $\chi(c)=-1$ (respectively, $\chi(c)=1$).
Lemma 2.19.
Given an $E$-rational Serre compatible system $\{\rho_{\lambda}\}$ and one extends the scalars of each $\rho_{\lambda}$ to $\overline{E}_{\lambda}$. If some $\rho_{\lambda_{0}}$ satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi_{\lambda_{0}}$. Then each $\rho_{\lambda}$ satisfies $\rho_{\lambda}\cong\rho_{\lambda}^{\vee}\otimes\chi_{\lambda}$ for some character $\chi_{\lambda}$. Moreover if the simititude character $\chi_{\lambda_{0}}$ is totally odd (resp. totally even), then one can choose each simititude character $\chi_{\lambda}$ to be totally odd (resp. totally even).
Proof.
One has $\rho_{\lambda_{0}}\otimes\rho_{\lambda_{0}}^{\vee}\otimes\chi_{\lambda}\cong\rho_{\lambda_{0}}\otimes\rho_{\lambda_{0}}$. As $\rho_{\lambda_{0}}\otimes\rho_{\lambda_{0}}^{\vee}$ has a trivial subrepresentation, $\chi_{\lambda_{0}}$ is a subrepresentation of $\rho_{\lambda_{0}}\otimes\rho_{\lambda_{0}}$. By [BH24, Theorem 1.1], this $\chi_{\lambda_{0}}$ is locally algebraic. Hence after possibly enlarging $E$, it extends to a compatible system $\{\chi_{\lambda}\}$. Then by semisimplicity, $\rho_{\lambda}\cong\rho_{\lambda}^{\vee}\otimes\chi_{\lambda}$. Finally, as $\chi_{\lambda}$ can be written as a product of a finite image character and some power of cyclotomic character, $\chi_{\lambda}(c)$ is independent of $\lambda$ for a fixed complex conjugation $c$. ∎
One generates total oddness of a Galois character to higher dimension as following:
Definition 2.20.
Let $K$ be a totally real field and $E$ be a number field.
  1. (i)
    A $\lambda$-adic Galois representation $\rho:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})$ is called essentially self-dual and totally odd, if it either factors through $\text{GSp}_{n}$ with totally odd similitude character or it factors through $\text{GO}_{n}$ with totally even similitude character.
  2. (ii)
    An $E$-rational strictly compatible system $\{\rho_{\lambda}\}$ of $K$ is called essentially self-dual and totally odd, if each $\rho_{\lambda}$ is essentially self-dual and totally odd.
2.5. Automorphic Galois representations
A general way to show a Galois representation fits into a strictly compatible system is to show it is automorphic.
Theorem 2.21.
[BLGGT14, Theorem C]. Suppose $K$ is a totally real field. Let $n$ be an integer and $\ell\geq 2(n+1)$ be a prime. Let
$\rho:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{\mathbb{Q}_{\ell}})$
be a continuous representation. Suppose that the following conditions are satisfies.
  1. (i)
    (Unramified almost everywhere) $\rho$ is unramified at all but finitely many primes.
  2. (ii)
    (Odd essential self-duality) Either $\rho$ maps to $\text{GSp}_{n}$ with totally odd similitude character or it maps to $\text{GO}_{n}$ with totally even similitude character.
  3. (iii)
    (Potential diagonalizability and regularity) $\rho$ is potentially diagonalizable (and hence potentially crystalline) at each prime $v$ of $K$ above $\ell$ and for each $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ it has $n$ distinct $\tau$-Hodge-Tate weights.
  4. (iv)
    (Irreducibility) $\rho|_{\mathrm{Gal}_{K(\zeta_{\ell})}}$ is residually irreducible.
Then we can find a finite Galois totally real extension $K^{\prime}/K$ such that $\rho|_{\mathrm{Gal}_{K^{\prime}}}$ is automorphic. Moreover $\rho$ is part of a strictly pure compatible system of $K$.
The following result shows for 2-dimensional representations of totally real fields, one can rule out even cases when $\lambda$ is sufficiently large.
Proposition 2.22.
[CG13, Proposition 2.5]. Suppose $K$ is a totally real field. Let $\ell>7$ be a prime. Suppose
$\rho:\mathrm{Gal}_{K}\to\text{GL}_{2}(\overline{\mathbb{Q}_{\ell}})$
is a continuous representation. Assume:
  1. (i)
    $\rho$ is unramified outside of finitely many primes.
  2. (ii)
    $\text{Sym}^{2}\overline{\rho}|_{\mathrm{Gal}_{K(\zeta_{\ell})}}$ is irreducible.
  3. (iii)
    $\ell$ is unramified in $K$.
  4. (iv)
    For each place $v\mid\ell$ of $K$ and each $\tau:K_{v}\hookrightarrow\overline{\mathbb{Q}_{\ell}}$, the $\tau$-Hodge-Tate weights of $\rho|_{\mathrm{Gal}_{K_{v}}}$ is a set of 2 distinct integers whose difference is less than $(\ell-1)/2$ and $\rho|_{\mathrm{Gal}_{K_{v}}}$ is crystalline.
Then the pair $(\rho,\det\rho)$ is essentially self-dual and odd.
In particular, above results shows certain low dimensional subrepresentations of strictly compatible systems fit into strictly compatible systems.
Proposition 2.23.
[Hu23a, Proposition 2.12], [DWW24, Proposition 2.21]. Given an $E$-rational strictly compatible system $\{\rho_{\lambda}\}$ of some totally real field, then for all but finitely many $\lambda$,
  1. (i)
    If $\sigma$ is a 2-dimensional irreducible regular subrepresentation of $\rho_{\lambda}$, then $\sigma$ can be extended to a 2-dimensional regular irreducible strictly compatible system.
  2. (ii)
    If $\sigma$ is a 3-dimensional irreducible regular essentially self-dual subrepresentation of $\rho_{\lambda}$, then $\sigma$ can be extended to a 3-dimensional regular irreducible strictly compatible system.
One immediate consequence is the following.
Corollary 2.24.
Given an $E$-rational strictly compatible system $\{\rho_{\lambda}\}$ of some totally real field. If there exists infinitely many $\lambda$ such that the irreducible decomposition of $\rho_{\lambda}$
$\rho_{\lambda}=\oplus_{i}W_{i}$
consists of only 1 or 2 dimensional regular irreducible conponents. Then the compatible system $\{\rho_{\lambda}\}$ can be written as direct sum of some 1 or 2 dimensional irreducible strictly compatible systems.
2.6. Semisimple reduction
We extract the following results from the proof of [Hu23b, Theorem 1.4].
Proposition 2.25.
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{n}(\overline{E_{\lambda}})\}$ be an $E$-rational strictly compatible system of $\mathbb{Q}$. Consider its modulo $\lambda$ compatible system $\{\overline{\rho_{\lambda}}^{\text{ss}}\}$ by taking semisimple reductions.
  1. (i)
    Suppose for infinitely many $\lambda$ one has a 2-dimensional odd irreducible subrepresentation
    $\sigma_{\lambda}\subseteq\overline{\rho_{\lambda}}^{\text{ss}}$
    Then after possibly twisting each one of them by a fixed power of cyclotomic character, there are infinitely many such $\sigma_{\lambda}$ which are attached to a fixed cuspidal eigenform (in the sense of Serre modularity conjecture).
  2. (ii)
    Suppose for infinitely many $\lambda$ one has a 4-dimensional irreducible subrepresentation which is the tensor product of two 2-dimensional representation
    $\sigma_{\lambda}\otimes\sigma_{\lambda}^{\prime}\subseteq\overline{\rho_{\lambda}}^{\text{ss}}$
    such that $\sigma_{\lambda}$ is odd. Then after possible twist $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ with inverse power of cyclotomic characters, there are infinitely many such $\sigma_{\lambda}$ which are attached to a fixed cuspidal eigenform (in the sense of Serre modularity conjecture).
Sketch.
The idea is to bound weights and levels of the modular liftings of $\sigma_{\lambda}$ when $\lambda$ varies. Then the results follow from the fact there can only be finitely many eigenforms with fixed weight and level.
To bound weights in both statements one uses Fontaine-Laffaille theory, which relays on 2.1.(iii). To bound levels in (i) one makes use of 2.1.(iv). To bound levels in (ii), by a result of Tate [Se77, Theorem 5], it is enough to lift $\sigma_{\lambda}$ of decomposition subgroup $D_{p}$ to $\text{GL}_{2}$ with only finitely many possible levels, but only for those (finitely many) ramified prime $p$ belongs to the exceptional set $S$ of the compatible system . After fixing an isomorphism of fields $\mathbb{C}\cong\overline{\mathbb{Q}_{\ell}}$, the image of modular lifting of $\sigma_{\lambda}$ gives a finite solvable subgroup of $\text{PGL}(\mathbb{C})$. Then for each such (solvable hence rule out type $S_{5}$) type finite subgroup, the desired lifting is constructed in the proof. ∎
We use the following result in the proof. See also [PSW18, Corollary 1.3] to remove regularity condition.
Proposition 2.26.
[BLGGT14, Proposition 5.3.2] Given a regular $E$-rational strictly compatible system $\{\rho_{\lambda}\}$ of some number field $K$. There is a Dirichlet density 1 many rational primes $\ell$ such that for any $\lambda|\ell$ and any irreducible subrepresentation $\sigma\subseteq\rho_{\lambda}$, $\sigma|_{\mathrm{Gal}_{K(\zeta_{\ell})}}$ is residually irreducible.
3. The proof
3.1. The non Lie-irreducible case
If $\rho_{\lambda_{0}}$ is not Lie-irreducible. As the Hodge-Tate weights are distinct, by 2.14 one can write
$\rho_{\lambda_{0}}=\text{Ind}^{\mathbb{Q}}_{K}\sigma$
for some number field $K$ and Lie-irreducible representation $\sigma$ of $K$. There are three cases to consider.
3.1.1. $[K:\mathbb{Q}]=6$
In such case $\sigma$ is an algebraic $\lambda$-adic character. By class field theory after enlarging $E$, this $\sigma$ extends to an $E$-rational compatible system $\{\sigma_{\lambda}\}$. Then $\{\text{Ind}^{\mathbb{Q}}_{K}\sigma_{\lambda}\}$ also is an $E$-rational compatible system. By semisimplicity each $\rho_{\lambda}$ is isomorphic to $\text{Ind}^{\mathbb{Q}}_{K}\sigma_{\lambda}$. Then the irreducibility of all $\rho_{\lambda}$ comes from Mackey’s irreducibility criterion and regularity. In particular this $\{\rho_{\lambda}\}$ is induced by a one-dimensional strictly compatible system of $F$.
3.1.2. $[K:\mathbb{Q}]=3$
Suppose $\rho_{\lambda}$ is reducible. Since $\rho_{\lambda}$ and $\rho_{\lambda_{0}}$ have same formal bi-character by 2.9.(ii), the irreducible decomposition of such $\rho_{\lambda}$ either has the dimensional type $2+2+2$ or $2+4$. Here for example type $2+2+2$ means the irreducible decomposition of $\rho_{\lambda}$ is
$\rho_{\lambda}=W_{\lambda}\oplus W_{\lambda}^{\prime}\oplus W_{\lambda}^{\prime\prime}$
with $\dim W_{\lambda}=\dim W_{\lambda}^{\prime}=\dim W_{\lambda}^{\prime\prime}=2$. Recall $d(\rho_{\lambda})$ is the Dirichlet density of rational primes which have a split factor in at least one field appears in the Lie-irreducible decomposition of $\rho_{\lambda}$. For our $\rho_{\lambda_{0}}$:
Lemma 3.1.
$d(\rho_{\lambda_{0}})=\left\{\begin{array}[]{cc}1/3&\text{if}\ K/\mathbb{Q}\ \text{is Galois}\\ 2/3&\text{if}\ K/\mathbb{Q}\ \text{is non-Galois}\end{array}\right.$
Proof.
This is a standard application of Chebotarev’s density theorem. If $K/\mathbb{Q}$ is not Galois, denote by $L$ the Galois closure. Then $\text{Gal}(L/\mathbb{Q})=S_{3}$, the symmetric group of three elements $\{1,2,3\}$. We only need to consider unramified rational prime $p$ of $K/\mathbb{Q}$. It has a split factor in $K$ if and only if $p$ is completely split or $p\mathscr{O}_{K}=\mathfrak{P}_{1}\mathfrak{P}_{2}$ with inertia degrees of $\mathfrak{P}_{1},\mathfrak{P}_{2}$ to be 1 and 2. This equals to require either $p\mathscr{O}_{L}$ is completely split or $p\mathscr{O}_{L}=\mathfrak{P}_{1}\mathfrak{P}_{2}\mathfrak{P}_{3}$ with inertia degree of each $\mathfrak{P}_{i}$ is 2. Hence equals to require $\text{Frob}_{p}\in S_{3}$ is either identity or a transposition. By Chebotarev’s density theorem, we have:
$d(\rho_{\lambda_{0}})=\frac{\#\{e,(1,2),(1,3),(2,3)\}}{\#S_{3}}=2/3$
2.16 rules out the situation that some irreducible component of $\rho_{\lambda}$ is Lie-irreducible. We also assume we are not in the case that each irreducible component of $\rho_{\lambda}$ is induced by some character, since otherwise by regularity and class field theory, we can use the same method as in 3.1.1 to show $\rho_{\lambda_{0}}$ is reducible. The only remaining case is:
$\rho_{\lambda}=\text{Ind}^{\mathbb{Q}}_{F_{1}}\chi\oplus\text{Ind}^{\mathbb{Q}}_{F_{2}}\sigma$
where $F_{1}$ and $F_{2}$ are quadratic number field, and $\chi$ is a character of $F_{1}$, and $\sigma$ is 2-dimensional of $F_{2}$. But one calculates the following hence contradicts with 2.15 again.
Lemma 3.2.
$d(\rho_{\lambda})=\left\{\begin{array}[]{cc}1/4&F_{1}\not=F_{2}\\ 1/2&F_{1}=F_{2}\end{array}\right.$
Proof.
A rational prime $p$ has a split factor in both $F_{i}$ if and only if $p$ is completely split in both $F_{i}$, if and only if $p$ is completely split in $F_{1}F_{2}$. Hence by Chebotarev’s density theorem we have $d(\rho_{\lambda})=1/[F_{1}F_{2}:\mathbb{Q}]$. ∎
Remark 1.
We have shown each $\rho_{\lambda}$ is induced by some 2-dimensional $\sigma_{\lambda}$ of some cubic field $K_{\lambda}$. If $K/\mathbb{Q}$ is Galois, then all $K_{\lambda}$ coincide. This is because when $K/\mathbb{Q}$ is Galois, restricting $\rho_{\lambda_{0}}$ to $K$ has dimensional type $2+2+2$ since the Galois group acts transitively on the set of irreducible subrepresentations. But if some $K_{\lambda}\not=K$, restricting $\mathrm{Ind}_{K_{\lambda}}^{\mathbb{Q}}\sigma_{\lambda}$ to $K$ still is irreducible and induced by a 2-dimensional Lie-irreducible representation of $KK_{\lambda}$. This contradicts with 2.15. Moreover, as $K$ is also totally real, restrict $\{\rho_{\lambda}\}$ to $K$ then apply 2.23.(i), one obtains some $\sigma_{\lambda}$ fits into a compatible system. Hence by semisimplicity, one writes $\{\rho_{\lambda}\}$ as the induced representations of a 2-dimensional Lie-irreducible strictly compatible system of $K$.
3.1.3. $[K:\mathbb{Q}]=2$
In this case $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}\otimes\chi$ where $\chi$ is the nontrivial character of $\mathbb{Z}/2\mathbb{Z}$ induced by $K$. By semisimplicity this extends to an isomorphism of compatible systems $\rho_{\lambda}\cong\rho_{\lambda}\otimes\chi$.
Suppose $\rho_{\lambda}$ is reducible for infinitely many $\lambda$. By 2.24, one can assume each such $\rho_{\lambda}$ contains a dimension $\geq 3$ irreducible component $W_{\lambda}$. Suppose first $\dim W_{\lambda}=3$, then the irreducible decomposition of $\rho_{\lambda}$ has to be
$\rho_{\lambda}=W_{\lambda}\oplus W_{\lambda}^{\prime}$
with $\dim W_{\lambda}^{\prime}=3$ since otherwise would contradict with 2.15. If $W_{\lambda}\otimes\chi\cong W_{\lambda}^{\prime}$, since $\chi$ has finite image, after restricting to a finite extension, the two components $W_{\lambda}$ and $W_{\lambda}^{\prime}$ are isomorphic contradicts with regularity of the compatible system. Hence one has $W_{\lambda}\cong W_{\lambda}\otimes\chi$. Twist $\rho_{\lambda}$ with a suitable power of cyclotomic character, one can assume $W_{\lambda}$ has algebraic monodromy group $\text{G}_{W_{\lambda}}=\text{GO}_{3}$ or $\text{GL}_{3}$. But this would contradict with $W_{\lambda}\cong W_{\lambda}\otimes\chi$.
Assume $\dim W_{\lambda}=4$ or 5. By 2.16, one can assume each of the irreducible components of $\rho_{\lambda}$ cannot be Lie-irreducible. We also assume not every irreducible component of $\rho_{\lambda}$ is induced by some character since otherwise we use method in 3.1.1 to contradict with $\rho_{\lambda_{0}}$ being irreducible. Then the only remaining case is:
$\rho_{\lambda}=\text{Ind}^{\mathbb{Q}}_{F}\sigma_{\lambda}\oplus\text{Ind}^{\mathbb{Q}}_{F}\chi_{\lambda}$
where $\sigma_{\lambda}$ is 2-dimensional Lie-irreducible of $F$, and $\chi_{\lambda}$ is a character of $F$, and $F$ is a quadratic number field. For $\lambda$ sufficiently large, the 2-dimensional part $W_{\lambda}^{\prime}$ fits into a strictly compatible system by 2.23.(i). We will show the four-dimensional part is potential automorphic. Then the semisimplicity of compatible system would contradict the irreducibility of $\rho_{\lambda_{0}}$.
It is enough to check the conditions in 2.21. (i) is obvious, (iii) can be obtained from [BLGGT14, Lemma 1.4.3.(2)]. To check (ii), as $\sigma_{\lambda}$ is Lie-irreducible, the algebraic monodromy group of $W_{\lambda}$ must contain $\text{SL}_{2}\times\text{SL}_{2}$. Since $\rho_{\lambda}$ satisfies $\rho_{\lambda}\cong\rho_{\lambda}^{\vee}\otimes\chi_{\lambda}$, by dimensional reason one has $W_{\lambda}\cong W_{\lambda}^{\vee}\otimes\chi_{\lambda}$. Hence the algebraic monodromy group of $W_{\lambda}$ must factor through $\text{GSp}_{4}$. Moreover, one can always twist $\chi$ to $\{\chi_{\lambda}\}$ making them odd. Hence $W_{\lambda}$ is essentially self-dual and totally odd. Finally, we check (iv), namely for $\lambda$ suffciently large, $\tau=(\text{Ind}_{F}^{\mathbb{Q}}\sigma_{\lambda})|_{\mathbb{Q}(\zeta_{\ell})}$ is residually irreducible, where $\ell$ is the unique rational prime divided by $\lambda$.
For $\ell$ sufficiently large, $\mathbb{Q}(\zeta_{\ell})$ and $F$ are linearly disjoint. Hence one has:
$\text{Res}_{\mathbb{Q}(\zeta_{\ell})}\text{Ind}_{F}^{\mathbb{Q}}\sigma_{\lambda}=\text{Ind}_{F(\zeta_{\ell})}^{\mathbb{Q}(\zeta_{\ell})}\text{Res}_{F(\zeta_{\ell})}\sigma_{\lambda}$
As $\sigma_{\lambda}$ is Lie-irreducible, $\text{Res}_{F(\zeta_{\ell})}\sigma_{\lambda}$ is irreducible. But this restriction is of type A since it has dimension 3, hence the induced representation also is of type A. Hence by 2.11.(iv) it is residually irreducible.
Remark 2.
By same method as in 1, $\{K_{\lambda}\}$ coincide for all but finitely many $\lambda$.
Remark 3.
In this non Lie-irreducible case, our proof also works for compatible systems over totally real field, except that to apply potential automorphy theorem in 3.1.3, one has to assume for all complex conjugation $c$, the similitude character $\chi_{\lambda_{0}}(c)$ has same sign. Note that 1 also is true for totally real field, and under above extral condition 2 also is true for totally real fields. We summarize the result as following.
Theorem 3.3.
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational regular strictly compatible system of $\mathbb{Q}$. Suppose some $\rho_{\lambda_{0}}$ is irreducible but not Lie-irreducible. Suppose further that $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ and the sign $\chi(c)$ is independent of complex conjugation $c$, then $\rho_{\lambda}$ is irreducible for all but finitely many $\lambda$.
3.2. The Lie-irreducible case
Given a Lie-irreducible representation $\rho$, denote by $\rho^{\text{der}}$ the restriction to the derived subgroup, which also is irreducible. In the proof we will call the unique $\mathbb{Z}$-model of $\rho^{\text{der}}$ the semisimple type of $\rho$. For 6-dimensional Lie-irreducible representation there are 7 semisimple types:
  1. (a)
    $(\text{SL}_{2},\text{Sym}^{5}(\text{std}))$
  2. (b)
    $(\text{SL}_{2}\times\text{SL}_{2},\text{std}\otimes\text{Sym}^{2}(\text{std}))$
  3. (c)
    $(\text{SO}_{6},\text{std})$
  4. (d)
    $(\text{Sp}_{6},\text{std})$
  5. (e)
    $(\text{SL}_{3},\text{Sym}^{2}(\text{std}))$
  6. (f)
    $(\text{SL}_{2}\times\text{SL}_{3},\text{std}\otimes\text{std})$
  7. (g)
    $(\text{SL}_{6},\text{std})$
Here symbol $\text{Sym}^{r}(\sigma)$ denotes the $r$th symmetric power of a representation $\sigma$, and Std stands for the standard representation. We prove the following result:
Proposition 3.4.
Given a 6-dimensional $E$-rational regular strictly compatible system $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$. Assume one $\rho_{\lambda_{0}}$ is Lie-irreducible.
  1. (i)
    Assume $\rho_{\lambda_{0}}$ has semisimple type (a),(e),(f),(g), then each $\rho_{\lambda}$ is Lie-irreducible with same semisimple type.
  2. (ii)
    Assume $\rho_{\lambda_{0}}$ has semisimple type (b), then each $\rho_{\lambda}$ is Lie-irreducible with same semisimple type for all but finitely many $\lambda$.
  3. (iii)
    Assume $\rho_{\lambda_{0}}$ is fully polarisable (this means (c) or (d)), then each $\rho_{\lambda}$ is Lie-irreducible with same semisimple type for all but finitely many $\lambda$.
The proof of (i) is essentially the same as in [Hu23b]. Since the conditions here are different and also for completeness, we summarize the proof here.
3.2.1. Cases (a) and (g)
In case (a) the formal character of $\textbf{G}_{\lambda_{0}}^{\text{der}}$ is $\{x^{-5},x^{-3},x^{-1},x,x^{3},x^{5}\}$ and cannot be decomposed as disjoint union of two formal characters of lower dimensional irreducible representations. In case (g) the derived subgroup $\textbf{G}_{\lambda_{0}}^{\text{der}}=\text{SL}_{6}$ has maximal rank 5 forces each $\textbf{G}_{\lambda}^{\text{der}}=\text{SL}_{6}$. Hence in both cases all $\rho_{\lambda}$ are irreducible with same semisimple type.
3.2.2. Cases (e) and (f)
The formal character of $\textbf{G}_{\lambda_{0}}^{\text{der}}$ in case (e) is $\{x^{2},xy,y^{-1},y^{2},x^{-1},x^{-2}y^{-2}\}$, in case (f) is $\{xy,xz,xy^{-1}z^{-1},x^{-1}y,x^{-1}z,x^{-1}y^{-1}z^{-1}\}$. They both do not contain zero weight and pairwise inverse weights. Hence if some $\rho_{\lambda}$ is reducible, it cannot have 1 or 2-dimensional components. Hence the only possible reducible $\rho_{\lambda}$ has dimensional type 3+3. But since there cannot be any zero weights in formal characters in both cases, the two 3-dimensional conponents must have semisimple type $\text{SL}_{3}$. Then in particular the semisimple rank is 4, which contradicts both cases. Hence every $\rho_{\lambda}$ is irreducible with same semisimple type.
3.2.3. Cases (c) and (d)
Since the formal character of $\textbf{G}_{\lambda_{0}}^{\text{der}}$ in both cases is $\{x,x^{-1},y,y^{-1},z,z^{-1}\}$, if there are infinitely many $\rho_{\lambda}$ which is reducible, it must contains a 2-dimensional irreducible subrepresentation. By 2.23.(i) for some $\lambda_{1}$, this 2-dimensional irreducible subrepresentation $\sigma_{\lambda_{1}}$ would fit into a compatible system $\{\sigma_{\lambda}\}$. Define a new strictly compatible system as:
$\{\rho_{\lambda}\oplus\sigma_{\lambda}\}$
By 2.7, $\rho_{\lambda_{1}}\oplus\sigma_{\lambda_{1}}$ has the same algebraic monodromy group as $\rho_{\lambda_{1}}$, hence has semisimple rank 3. However by Goursat’s lemma, the derived subgroup of algebraic monodromy group of $\rho_{\lambda_{0}}\oplus\sigma_{\lambda_{0}}$ is either $\text{Sp}_{6}\times\text{SL}_{2}$ or $\text{SO}_{6}\times\text{SL}_{2}$, which has semisimple rank 4. This contradicts with 2.9(ii).
Next we distinguish the semisimple type $\text{Sp}_{6}$ and $\text{SO}_{6}$. Assume our compatible system $\{\rho_{\lambda}\}$ have infinitely many terms with semisimple type $\text{SO}_{6}$ and infinitely many terms with semisimple type $\text{Sp}_{6}$. If the similitude character $\chi_{\lambda_{0}}$ is odd (resp. even), one can choose all similitude characters $\chi_{\lambda}$ of $\rho_{\lambda}$ to be odd (resp. even). Assume $\rho_{\lambda}$ belongs to either one of the following infinite sets:
  1. (i)
    In the odd similitude characters case, the set of $\rho_{\lambda}$ with semisimple type $\text{Sp}_{6}$.
  2. (ii)
    In the even similitude characters case, the set of $\rho_{\lambda}$ with semisimple type $\text{SO}_{6}$.
We will show in both cases some $\rho_{\lambda}$ is potentially automorphic. Then the result follows since a strictly compatible system (of a totally real field) attached to a regular algebraic, essentially self-dual, cuspidal automorphic representation has the desired property that either all of the representations factor through $\text{GSp}_{n}$ or all of them factor through $\text{GO}_{n}$.
We only need to check conditions in 2.21. And it is only condition (iv) requires further explaination. We first show $\overline{\rho_{\lambda}}^{\text{ss}}$ is irreducible for all but finitely many $\lambda$. This also is partial proof of 1.3.(iii). Note if the semisimple type is $\text{SO}_{6}$ which is type A, by 2.11.(iii) $\overline{\rho}_{\lambda}^{\text{ss}}$ is irreducible when restricting to $\mathbb{Q}(\zeta_{\ell})$ for all but finitely many $\lambda$.
So we are in case (i) above. As the formal bi-character of $\rho_{\lambda}$ and $\overline{\rho}_{\lambda}^{\text{ss}}$ coincide. If the latter is reducible, there exists a 2-dimensional irreducible component $\overline{\sigma}_{\lambda}$. Oddness and regularity guaratee this $\sigma_{\lambda}$ is odd. Hence by 2.25.(i), there exists a 2-dimensional strictly compatible system $\{\varphi_{\lambda}\}$ such that for infinitely many $\lambda$ one has $\overline{\varphi}_{\lambda}^{\text{ss}}\cong\overline{\sigma}_{\lambda}$. Construct an 8-dimensional strictly compatible system:
$\{\psi_{\lambda}:=\rho_{\lambda}\oplus\varphi_{\lambda}\}$
Then on the one hand by Goursat’s lemma $\textbf{G}_{\lambda}^{\text{der}}=\text{Sp}_{6}\times\text{SL}_{2}$ which has rank 4. On the other hand for some $\lambda$ one has $\overline{\psi}_{\lambda}^{\text{ss}}=\overline{\rho}_{\lambda}^{\text{ss}}\oplus\overline{\sigma}_{\lambda}$ which by 2.7 has rank 3. This is a contradiction.
Next we show $\overline{\rho}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If on the contrary this is reducible, by [CG13, Lemma 4.3] $\overline{\rho}_{\lambda}^{\text{ss}}$ is of induced type. Hence the Dirichlet density $d(\overline{\rho}_{\lambda}^{\text{ss}})$ is strictly smaller than 1. However, by [Pa19, Proposition 3.4.9.(i)] this density is the Dirichlet density of the set of primes whose Frobenius has zero trace. But since $\rho_{\lambda}$ is Lie-irreducible, this density should be 1, a contradiction.
3.2.4. Case (b)
Since the formal bi-character is $\{x,x^{-1},xy,x^{-1}y,xy^{-1},x^{-1}y^{-1}\}$. If $\rho_{\lambda}$ is reducible, the irreducible decomposition has to be $\rho_{\lambda}=W_{\lambda}\oplus W_{\lambda}^{\prime}$ with $\dim W_{\lambda}=2,\dim W_{\lambda}^{\prime}=4$. As $\text{Sp}_{4}$ does not have 2-dimensional irreducible representations, $\textbf{G}_{W_{\lambda}^{\prime}}^{\text{der}}$ has to be $\text{SO}_{4}$. Hence the representation factors through $\text{GO}_{4}$.
By [LY16, Section 3.3] (more generally see also [DWW24, Theorem 2.13]), one can write those infinitely many $W_{\lambda}^{\prime}$ as
(1) $W_{\lambda}^{\prime}=\sigma_{\lambda}\otimes\sigma_{\lambda}^{\prime}$
where $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ both are irreducible, 2-dimensional, unramified at all but finitely many rational primes and crystalline when restrict to $\mathrm{Gal}_{\mathbb{Q}_{\ell}}$, where $\ell$ is the unique rational prime divided by $\lambda$.
We show for $\lambda$ sufficiently large, both $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ fit into compatible systems. Then by semisimplicity this contradicts with $\rho_{\lambda_{0}}$ being irreducible.
It is enough to assume both $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ are Lie-irreducible. We check the conditions in 2.21. (i) is obvious, (iii) as before follows from [BLGGT14, Lemma 1.4.3.(2)]. (iv) follows from 2.11.(iii). We use 2.22 to check the oddness condition (ii). In those four conditions, (i) and (iii) are obvious. (ii) follows from 2.11.(iii). Finally for the boundness of difference of Hodge-Tate weights, this comes from the facts that Hodge-Tate weights of all $\rho_{\lambda}$ are the same, and the set of differences of Hodge-Tate weights of $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ are contained in the set of differences of Hodge-Tate weights of $\rho_{\lambda}$.
Remark 4.
In our proof of 1.3(i), everything works for a regular strictly compatible system of a general totally real field $F$, except for this case (b). Here we need to write a Galois representation $\text{Gal}_{F}\to\text{GO}_{4}(\overline{E_{\lambda}})$ as the tensor product of two 2-dimensional representations with certain $\ell$-adic hodge theoretical properties, which can be done for $F=\mathbb{Q}$. If however the strictly compatible system of a totally real field is pure, then by [Pa19, Section 3.2], one still have (1) with the desired properties. Then our proof works fine for such compatible systems.
3.3. Residual irreducibility
We summarize the proof of 1.3.(iii). The non Lie-irreducible cases can be proven the same as in the end of 3.1.3. For the Lie-irreducible cases, type A ones follow from 2.11.(iii); the fully symplectic case can be proven the same as in 3.2.3.
3.4. Proof of 1.4
As we are assuming purity of the compatible system, 4 asserts 1.3(i) is true for compatible systems of totally real fields. Hence we can assume each $\rho_{\lambda}$ is reducible. We first prove two lemmas:
Lemma 3.5.
For all but finitely many $\lambda$, the irreducible component of $\rho_{\lambda}$ which either has dimension 1,2,4 or has dimension 3 with semisimple type $\text{SO}_{3}$, fits into a strictly compatible system.
Proof.
By class field theory and 2.23, for all but finitely many $\lambda$, if some irreducible component $\sigma_{\lambda}$ of $\rho_{\lambda}$ either has dimension $\leq 2$ or semisimple type $\text{SO}_{3}$, then it fits into a compatible system.
Assume $\dim\sigma_{\lambda}=4$. This is essentially self-dual and totally odd due to dimensional resason. Hence to apply 2.21 we only need to check $\overline{\sigma}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If it is induced, then the same method as in 3.1.3 shows $\overline{\sigma}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If it is Lie-irreducible, if the semisimple type is of type A, then 2.11.(iii) shows the desired result. Hence we only need to assume the semisimple type of $\sigma_{\lambda}$ is fully symplectic. Replace $\text{Sp}_{6}$ with $\text{Sp}_{4}$, the same proof as in 3.2.3 applies. ∎
Lemma 3.6.
Under the purity condition, there cannot be some $\rho_{\lambda_{0}}$ which contains a 3-dimensional irreducible component $W_{\lambda_{0}}$ with semisimple type $\text{SL}_{3}$.
Proof.
Suppose on the contrary such $W_{\lambda_{0}}$ exists. By essential self-duality, one must have irreducible decomposition $\rho_{\lambda_{0}}=W_{\lambda_{0}}\oplus W^{\prime}_{\lambda_{0}}$ with $W^{\prime}_{\lambda_{0}}=W^{\vee}_{\lambda_{0}}\otimes\chi_{\lambda_{0}}$. Hence there cannot be infinitely many $\lambda$ such that the irreducible components of $\rho_{\lambda}$ are all of types mentioned in 3.5. (If there are infinitely many $\lambda$ such that $\rho_{\lambda}$ has dimensional type 3+3 and each component has semisimple type $\text{SO}_{3}$, then those components extends to a compatible system and contains $W_{\lambda_{0}}$. The semisimple ranks do not match).
Hence the only possible dimensional type of $\rho_{\lambda}$ other than 3+3 is 5+1. Denote by $d$ the Dirichlet density of $\lambda$ such that $\rho_{\lambda}$ has dimensional type 5+1. If $d>0$, then by 2.26 there are infinitely many those 5-dimensional component satisfies 2.21.(iv). Other conditions in 2.21 are easy to verify. Hence one writes $\{\rho_{\lambda}\}$ as a direct sum of a 5-dimensional compatible system and a 1-dimensional compatible system. This contradicts with the decomposition of $\rho_{\lambda_{0}}$. So we must have $d=0$. But then the Dirichlet density of $\lambda$ such that $\rho_{\lambda}$ contains a 3-dimensional component with semisimple type $\text{SL}_{3}$ would be 1. This contradicts with [PT15, Lemma 1.6]. ∎
To finish the proof, denote by $r$ the smallest integer such that there are infinitely many $\lambda$ such that each the irreducible components of $\rho_{\lambda}$ have dimenion $\leq r$. Then 3.5 asserts the result is true for $r\leq 4$. Assume $r=5$, then one again uses 2.26 and checks conditions in 2.21 to show the 5-dimensional components extend to a strictly compatible system. This finishes the proof.
Acknowledgement
I would like to express my deepest gratitude to my advisor, Chun Yin Hui, for his invaluable guidance, insightful suggestions, unwavering support, and encouragement throughout not only this research, but also my Ph.D. journey. I am deeply appreciative of his time, effort, and dedication. His mentorship has been a cornerstone of my academic journey.